In integers, 4x – 1 = 8 has                                solution.
 In integers, 4x – 1 = 8 has                                solution.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given integers, 4x–1 = 8
Transpose -1 to RHS then it becomes 1.
4x = 8 + 1
4x = 9
x = 9/4
In integers, 4x–1 = 8 has no solution.


Thanks